A manufacturer wishes to recieve $4800 for a batch of items. When 1200 items were found to be defective, the selling cost per item increased by $0.20. How many items were manufactured?
Answer by richwmiller(17219)

(x-1200)*.20=4800
.2x-.2*120 =4800
(x-1200)*.20=4800
x = 25200.